Let X, Y be independent and identically distributed random variables, each having Poisson distribution with rate parameter λ. Find the probability generating function of W = X + Y and hence the mean and variance of W


1
Expert's answer
2022-05-10T23:31:40-0400

"X" has a Poisson distribution. It means that "X" takes values from the set of nonnegative integers "0,1,2,..." and "P(X=k)=\frac{\lambda^ke^{-\lambda}}{k!}", where "k=0,1,2,...". For "Y" we have: "P(Y=n)=\frac{\lambda^ne^{-\lambda}}{n!}", "n=0,1,2,...". A random variable "W=X+Y" has a Poisson distribution with parameter "2\lambda". It is a well-known fact. It can be also checked directly: "P(W=n)=P(X+Y=n)=\sum_{k=0}^nP(X=k,Y=n-k)=\sum_{k=0}^n\frac{\lambda^ke^{-\lambda}}{k!}\frac{\lambda^{n-k}e^{-\lambda}}{(n-k)!}=\sum_{k=0}^n\frac{\lambda^ne^{-2\lambda}}{k!(n-k)!}=\frac{e^{-2\lambda}}{n!}\sum_{k=0}^n\frac{n!\lambda^n}{k!(n-k)!}=\frac{(2\lambda)^n}{n!}e^{-2\lambda}"

We used a binomial formula. Remind that the probability generating function for "W" has the form:

"G(z)=E[z^W]=\sum_{x=0}^{+\infty}p(x)z^x=\sum_{x=0}^{+\infty}\frac{(2\lambda)^x}{x!}e^{-2\lambda}z^x=e^{-2\lambda}\sum_{x=0}^{+\infty}\frac{(2\lambda z)^x}{x!}=e^{2\lambda(z-1)}"

We used a Taylor series expansion for "e^{2\lambda z}". The latter is a known formula. Substitute "z=e^t". We receive: "M(t)=E[e^{tW}]=e^{2\lambda(e^t-1)}". This is a moment-generating function. Using the known formulae, we have: "E[W]=M'(0)=(2\lambda e^te^{2\lambda(e^t-1)})|_{t=0}=2\lambda".

"E[W^2]=M''(0)=(2\lambda e^te^{2\lambda(e^t-1)}+4\lambda^2 e^{2t}e^{2\lambda(e^t-1)})|_{t=0}=2\lambda+4\lambda^2".

"Var[W]=E[W^2]-(E[W])^2=2\lambda+4\lambda^2-4\lambda^2=2\lambda"

Answer: The probability generating function of "W" is: "G(z)=E[z^W]=e^{2\lambda(z-1)}". The mean is: "E[W]=2\lambda", "Var[W]=2\lambda".